OpenStudy (anonymous):

for a projectile launched upwards from level ground if the speed of a projectile at the maximum height is half the speed of projection u its horizontal range is

OpenStudy (anonymous):

clue v0x=v0/2

OpenStudy (anonymous):

at the maximum height, the vertical component of velocity becomes zero. so only the horizontal component (u costheta) is present. from the given info. , \[u \cos \theta = u/2\] \[\cos \theta = 1/2\] \[\theta = 60\]

OpenStudy (anonymous):

solve it on your own from here. and i'd rather suggest that you go through the theory part of the chapter before starting with the numerical problems

OpenStudy (anonymous):

Write your solution here if it is not correct, we will help you.

OpenStudy (anonymous):

i got it thanxxx guys